# Scene-Specific Trajectory Sets: Maximizing Representation in Motion Forecasting

### v1

@@ Abstract
Representing diverse and plausible future trajectories of actors is crucial for motion forecasting in autonomous driving. However, efficiently capturing the true trajectory distribution with a compact set is challenging. In this work, we propose a novel approach for generating scene-specific trajectory sets that better represent the diversity and admissibility of future actor behavior. Our method constructs multiple trajectory sets tailored to different scene contexts, such as intersections and non-intersections, by leveraging map information and actor dynamics. We introduce a deterministic goal sampling algorithm that identifies relevant map regions and generates trajectories conditioned on the scene layout. Furthermore, we empirically investigate various sampling strategies and set sizes to optimize the trade-off between coverage and diversity. Experiments on the Argoverse 2 dataset demonstrate that our scene-specific sets achieve higher plausibility while maintaining diversity compared to traditional single-set approaches. The proposed Recursive In-Distribution Subsampling (RIDS) method effectively condenses the representation space and outperforms metric-driven sampling in terms of trajectory admissibility. Our work highlights the benefits of scene-aware trajectory set generation for capturing the complex and heterogeneous nature of actor behavior in real-world driving scenarios.
@@
I Introduction
@@
Figure 1: Trajectory set, which represents the possible full actor representation within the scene
@@ III Problem Formulation
While promoting diversity ($D$) among the trajectories within $T^{\prime}$, it is important to note that coverage alone is insufficient to ensure accurate and plausible output. $T^{\prime}$ should also be realistic or plausible ($P$), and therefore should not contain states that are physically infeasible to reach for a given scene $S$ in $M_{i}$. Additionally, $S$ includes lane information that delineates the probable motion paths an actor can choose to reach a goal position. Within this work, a Focal Track ($FT$) is of an object type car.
@@
Figure 2: Candidate set for intersection scenes
Figure 3: Candidate set for non-intersection scenes
IV Scene-Specific Set Generation
Figure 4: Past trajectory is shown along with the Future trajectory of an FT. The Grey edges represent the connectivity between different lanes with blue nodes representing the source nodes and Violet denotes the possible goal lanes which the FT can move along
IV-A Algorithm
$L_{goal}=f(S,\theta)$ (4)
Given a vector map of $S$ and the state information of $FT$, the algorithm constructs a directed graph $G$ from the map. The last observation state ($s_{i}^{0}$) of the $FT$ is used to find the nearest nodes by filtering for edges with parallel direction vectors to that of the $FT$. This helps us find for the edges which have a higher similarity value w.r.t the $FT$’s travelling direction, mathematically given through the dot product between the two normalized vectors (thresholding it at an angle of 10). Within this pool of filtered edges, we choose a source node which fulfills the property of being closer to $FT$; found by calculating the L1-norm between $s_{i}^{0}$ and the nearest nodes.
A depth-first search (DFS) is performed on each of the nearest nodes to find the goal lanes. For each goal lane, the algorithm checks if it is in an intersection. If any one of the lanes $l$ in $L_{goal}$ is present in an intersection, the trajectory is returned with a scenario tag of 1; otherwise 0. This process allows for the categorization of scenes into two bins based on the presence of goal lanes in an intersection, as can be seen from algorithm 1.
@@ IV-A Algorithm
1:Vector-Map $M$, State Information of Focal Track $FT$
@@
4:     $S_{past}$, $\theta$leftarrow Query for the focal track, where $\theta$ is the heading angle
5:     $R\leftarrow$ Calculate Rotation Matrix($\theta$)
@@
7:     $n_{source}\leftarrow$ Find the source node using L1-norm between $S_{t}[-1]$ and $N_{nearest}$
@@
Running algorithm 1 across the training dataset yields two different candidate sets that reflect the local representation of the scene, as shown in fig. 3 and fig. 3. The trajectories corresponding to the scenes are normalized to $FT$’s last observation state using the rotation matrix $R$ from line 4 of algorithm 1. This approach maximizes the probability of finding trajectories closer to the ground truth without compromising on the shortcomings of earlier approaches which utilize a single set. To optimize the candidate sets further, we need to determine two key parameters: the optimal set size and the sampling strategy. The sampling approach plays a crucial role in reducing the candidate sets to an optimal size suitable for the intended application.
V Metrics
@@
$DAC=\frac{\kappa-n}{\kappa}$
where, $\kappa$ represents the given set size and $n$ denotes the number of non-compliant trajectories which goes off-road. A higher DAC score provides an indication of the set’s density in relation to the scene. We use the same methodology as described in [5] to get the DAC value per scene.
By integrating the DAC and Euclidean metrics, we enhance our assessment of the set’s quality, which hinges on its closeness to the ground truth trajectories. A higher DAC score signifies that more trajectories adhere to drivable areas, whereas lower minADE and minFDE values indicate closer alignment with the ground truth. Employing both diversity and scene compliance metrics ensures that the optimized set of trajectories maximizes the representativeness of the set, providing a comprehensive portrayal of the true population dynamics w.r.t the overall map geometry.
VI Experiments
@@
TABLE I: Comparison of coverage-based approach vs. random sampling from the training distribution.
@@
A trajectory set for a training distribution $T_{full}$ is constructed by parsing through each scene and querying for the focal track’s future position. Since each of the trajectory is normalized with their respective $R$, we get a set which resembles fig. 1.
VI-A Evaluation of a single set
Validating the effectiveness of the sets is crucial to comprehend the influence of two primary parameters which govern diversity and plausibility. We employ the metric-driven method [12] and a random sampling to demonstrate the efficacy of our proposed scene-specific set generation concept. Similar to the Covernet [3]; metric-driven method also employs a minADE driven bagging algorithm to reduce each of the candidate sets, we have acquired from the algorithm 1. Random sampling (Ran-Samp) is used as a baseline, and we report the mean values as it exhibits a variance of ±$\pm\text{2\%}$%.
We compute a theoretical Lower Bound (LB) which measures the quality of the constructed set validated against the trajectories from the training distribution. This lower bound is derived from the minADE between the ground truth trajectory and the closest trajectory within the set. The proposed approach enables us to assess the quality and diversity of the generated trajectory sets. By comparing the performance of our method against the baseline and analyzing the lower bounds, we gain insights into how effective our set construction is in capturing diverse and plausible trajectories. table I illustrates the effect of metric-driven sets when compared with randomly sampled sets generated from the complete training distribution.
As the set size increases, the LB of the metrics linearly decreases, this is attributed to the presence of more trajectories which lie closer to the GT trajectory, thereby providing good coverage. Sets generated through the metric-driven reduction method perform well as expected, since the cost function enforces diversity between the trajectories in the set when compared with Ran-Samp.
On the other hand, from fig. 5 emerges a wholly different picture. The normalized rate of surviving trajectories, that which are compliant with the scene, is substantially higher for the Ran-Samp and amounts to $10\%$ more trajectories irrespective of the set size.
@@ VI-A Evaluation of a single set
VI-B Evaluation of scene-specific sets
Utilizing algorithm 1, we generated two distinct candidate sets: one pertaining to intersections and the other to non-intersections. Subsequently, for each candidate set, adhering to their corresponding sampling criteria, we performed extensive analysis, as can be seen from the table II.
TABLE II: Quantitative comparison of Scene-specific sets with different methods on Argoverse training distribution
@@ VI-B Evaluation of scene-specific sets
Method Set Size LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ set ()($\downarrow$)( ↓ ) DAC()($\uparrow$)( ↑ ) LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ set ()($\downarrow$)( ↓ ) DAC()($\uparrow$)( ↑ )
1k 0.59 0.67 0.694 0.72 0.73 0.518
1.5k 0.53 0.62 0.673 0.66 0.67 0.500
2k 0.50 0.59 0.677 0.62 0.63 0.490
Metric Driven 2.5k 0.47 0.56 0.664 0.59 0.60 0.484
3k 0.44 0.54 0.704 0.57 0.58 0.477
4k 0.41 0.52 0.692 0.54 0.55 0.468
5k 0.38 0.49 0.687 0.51 0.52 0.469
1k 0.72 0.83 0.827 0.874 0.89 0.691
1.5k 0.64 0.74 0.831 0.789 0.81 0.664
2k 0.60 0.68 0.823 0.729 0.74 0.663
Ran-Samp 2.5k 0.56 0.65 0.842 0.690 0.70 0.678
3k 0.54 0.63 0.833 0.659 0.68 0.692
4k 0.49 0.59 0.835 0.606 0.62 0.681
5k 0.45 0.55 0.863 0.579 0.58 0.679
@@
The intuition behind reporting LB.minADE w/ set, is to establish a baseline to compare how a method (metric-driven, ran-samp, RIDS); pertaining to a specific set constructed from a full training distribution, would fare against different scenes (non-intersection/intersection). This approach allows for a fair comparison between different set construction methods and helps identify the most suitable method for specific types of scenes.
The DAC metric (normalized surviving trajectories per scene) exhibits relative stability across different set sizes for each method and scenario combination. The Ran-Samp method consistently shows higher DAC values compared to metric-driven, suggesting that it is more effective in generating a set where a larger proportion of the trajectories remain plausible. When comparing the two scenario types, it is evident that both methods yield poorer performance in intersection scenarios compared to non-intersections, as measured between corresponding LB.minADE w/ set and LB.minADE metrics. This outcome is anticipated due to the increased complexity and diversity of potential trajectories at intersections. However, the performance degradation from non-intersection to intersection scenarios is less evident for Ran-Samp than for metric-driven. As set sizes increase, the displacement error narrows, and the number of surviving trajectories experiences a significant increase, a trend that is also observable in the intersection sets.
We also created a new subsampling algorithm, RIDS or Recursive In Distribution Subsampling, which is a method where we remove duplicate trajectories from a set constructed using Ran-samp. If two trajectories share proximity to each other and lie within a threshold of 0.2m to each other, we consider this trajectory redundant and arbitrarily remove one of them. This leaves us with a set which is underweighted; to counter this, we perform subsampling from the training distribution and recursively perform this step of removing and adding trajectories until the desired $T^{\prime}$ is reached.
VII Results and Discussions
The ultimate goal of this work is to find the answer to the following question: what is the best conceptual framework that can maximize the probability of finding trajectories closer to the ground truth? This led us to investigate and propose a scene-specific set generation method, which, when viewed through the lenses of displacement errors and DAC, gives better performance compared to an uninformed set. The next step is to find the sampling method; from fig. 5, random sampling provides preliminary evidence, as it retains more trajectories than the metric-driven when one wants to achieve a dense set representation.
@@ VII Results and Discussions
Although the metric-driven method was able to provide guarantees for probabilistic completeness for maximizing coverage, it failed to capture the true distribution, providing insufficient exploration. On the other hand, sampling from a uniform distribution; RIDS was able to retain more trajectories and outperform the metric-driven method. The property of an efficient estimator is one that has the smallest possible variance, indicating that it is closest to the true value of the parameter being estimated. This reasoning that RIDS being an efficient estimator and functions effectively because in certain optimization situations, sampling from the underlying distribution can be more efficient than performing sampling based on some metrics.
Our experiments have shown that the scene-specific set generation method allows us to improve diversity, and more specifically, RIDS as a sampling method works relatively well when considering the environment, resulting in an improved DAC score. The survival rate for trajectories for sets constructed using Ran-Samp and RIDS outperforms metric-driven sampling methods, given that the minimal gain in displacement errors from metric-driven overshadows the presence of more plausible trajectories. A visual distinction is notable for the non-intersection set in fig. 3, which confirms with our prior, as most of the trajectories represent no significant yaw differences (going straight). The DAC score for the non-intersection with Ran-Samp depicts that cluster-based trajectory sampling works better than expected, since it can retain 86% of the trajectories for $T^{5k}_{set}$. This high score indicates that the set can handle the complex geometry of the real-world and is perfectly suited for the intended application. As we said at the start, this showcases that Scene-Specific Set generation can produce sets which can be tuned to represent the ground truth closer to real-world scenes than using a single set to represent diverse situations.
Looking at Ran-Samp set created with a threshold rate of 0.2 from table I; almost one could say that this imbalance between the set size is the reason for the improved DAC value. To counter that argument, we performed ablation studies to confirm that these improved DAC values remain improved when performing the RIDS sampling process for different threshold values.
VIII Conclusion

### v2

# Efficient Data Representation for Motion Forecasting: A Scene-Specific Trajectory Set Approach

@@ Abstract
Representing diverse and plausible future trajectories is critical for motion forecasting in autonomous driving. However, efficiently capturing these trajectories in a compact set remains challenging. This study introduces a novel approach for generating scene-specific trajectory sets tailored to different contexts, such as intersections and straight roads, by leveraging map information and actor dynamics. A deterministic goal sampling algorithm identifies relevant map regions, while our Recursive In-Distribution Subsampling (RIDS) method enhances trajectory plausibility by condensing redundant representations. Experiments on the Argoverse 2 dataset demonstrate that our method achieves up to a 10% improvement in Driving Area Compliance (DAC) compared to baseline methods while maintaining competitive displacement errors. Our work highlights the benefits of mining such scene-aware trajectory sets and how they could capture the complex and heterogeneous nature of actor behavior in real-world driving scenarios.
@@
I Prologue
The need for scene-specific trajectory sets arises from the inherent variability in driving environments. Different scenes—such as intersections, roundabouts, or straight roads—impose unique constraints on vehicle movement due to factors like road geometry, traffic rules, and potential interactions with other vehicles. A single set of trajectories cannot capture this diversity effectively because it lacks the contextual information needed to represent plausible behaviors across all scenarios. For example, at intersections, vehicles may turn left, right, or go straight, each requiring different trajectory predictions based on lane geometry and traffic signals. In contrast, on straight roads, fewer trajectory variations are needed because vehicles typically follow a more predictable path. By generating scene-specific trajectory sets, we can ensure that the predicted trajectories are not only diverse but also plausible within the specific context of each scene. Moreover, using a single set risks either under-representing possible outcomes (leading to unsafe predictions) or over-representing them (resulting in computational inefficiency). Scene-specific sets strike a balance by focusing computational resources on the most relevant trajectories for each scenario. This approach enhances both safety and performance in autonomous driving systems by ensuring that predictions align more closely with real-world constraints and behaviors.
Figure 1: A Set for (a) Intersection Scenes and (b) Non-Intersection Scenes
II Introduction
@@ IV Problem Formulation
While promoting diversity ($D$) among the trajectories within $T^{\prime}$, it is important to note that coverage alone is insufficient to ensure accurate and plausible output. $T^{\prime}$ should also be realistic and plausible ($P$), and therefore should not contain states that are physically infeasible to reach for a given scene $S$ in $M_{i}$. Additionally, $S$ includes lane information that delineates the probable motion paths an actor can choose to reach a goal position. Within this work, a Focal Track ($FT$) is of an object type car.
@@
V Scene-Specific Set Generation
@@
1:Vector-Map ($m$), State ($s_{i}$) Information of Focal Track $FT$
@@
4:     $s_{i}$, $\theta$leftarrow Query for the focal track states, where $\theta$ is the heading angle
5:     $R\leftarrow$ Calculate Rotation Matrix with $\theta$
@@
7:     $n_{source}\leftarrow$ Find the source node using L1-norm between $s_{i}^{0}$ and $N_{nearest}$
@@
V-A Algorithm
$L_{goal}=f(S,\theta)$ (4)
Given a vector map of $S$ and the state information of $FT$, the algorithm constructs a directed graph $G$ from the map. The last observation state ($s_{i}^{0}$) of the $FT$ is used to find the nearest nodes by filtering for edges with parallel direction vectors to that of the $FT$. This helps us find for the edges which have a higher affinity to the $FT$’s travelling direction, mathematically given through the dot product between the two normalized vectors (thresholding it at an angle of 10). Within this pool of filtered edges, we choose a source node which fulfills the property of being closer to $FT$; found by calculating the L1-norm between $s_{i}^{0}$ and the nearest nodes.
A Depth-First Search (DFS) is performed on each of the nearest nodes to find the goal lanes. For each goal lane, the algorithm checks if it is in an intersection. If any one of the lanes $l$ in $L_{goal}$ is present in an intersection, the trajectory is returned with a scenario tag of 1; otherwise 0. This process allows for the categorization of scenes into two bins based on the presence of goal lanes in an intersection. Running Algorithm 1 across the training dataset yields two clustered sets that reflect the local representation of the scene, as shown in Fig. 1(a) and Fig. 1(b). The trajectories corresponding to the scenes are normalized to $FT$’s last observation state using the rotation matrix $R$.
As we will see in the upcoming chapters, how this clustering maximizes the probability of finding trajectories closer to the ground truth without compromising on the shortcomings of earlier approaches which utilize a single set. To optimize the clustered sets further, we need to determine two key parameters: the optimal set size and the sampling strategy. The sampling approach plays a crucial role in reducing the clustered sets to an optimal size suitable for the intended application.
VI Metrics
@@
$DAC=\frac{|T^{\prime}|-n}{|T^{\prime}|}$
where, $|T^{\prime}|$ represents the given set size and $n$ denotes the number of non-compliant trajectories which goes off-road. A higher DAC score provides an indication of the set’s density in relation to the scene. We use the same methodology as described in [5] to get the DAC value per scene. By integrating the DAC and Euclidean metrics, we enhance our assessment of the set’s quality, which hinges on its closeness to the ground truth trajectories. A higher DAC score signifies that more trajectories adhere to drivable areas, whereas lower minADE values indicate closer alignment with the ground truth. Employing both diversity and scene compliance metrics ensures that the optimized set of trajectories maximizes the representativeness of the set, providing a comprehensive portrayal of the true population dynamics w.r.t the map geometry.
To evaluate the quality of our constructed trajectory set, we establish a theoretical Lower Bound (LB) through comparison with the $T_{full}$ which provides us with the required query trajectories  [^1] We use both query trajectory and ground truth trajectory interchangeably throughout this paper. The LB computation process operates as follows: for each query trajectory $q\in T_{full}$, we determine the closest matching trajectory $T^{*}$ for a given $T^{\prime}$ by calculating minADE, formally expressed as:
(a) Drivable Regions
(b) Node Construction for Traversable Lanes
(c) Reachable Lanes
(d) Complete Scene with Actor Tracks
Figure 2: Steps involved in clustering the scenes based on the road geometry and actor states are shown here. Past trajectory and Future trajectory of an FT are illustrated. In (a), Grey connections denote the connectivity between various lanes at an intersection. In (b), Blue edges depict the transformation of lane connections into node edges. Using Algorithm 1, the Source node, which corresponds to the closest node relative to the last observed state of the FT, is identified. Finally, the edges in (c) represent the plausible lanes the FT could potentially select.
$T^{*}=\operatorname*{arg\,min}_{q\in T_{full}}\text{ADE}(q,T^{\prime})$ (5)
The LB is then computed as the mean of these minimum displacement errors across all query trajectories:
[…]
This metric provides a quantitative measure of how well our constructed set captures the diversity and representativeness of the full trajectory distribution; with lower values indicating better coverage of the possible motion patterns. The LB. minADE w/ $T_{full}$ is an important measure for which the $q$ originates from their respective scene types (intersection or non-intersection), thereby representing ground truth (GT) samples from the scene-specific cluster. The constructed set, however, draws trajectories from $T_{full}$, enabling a comprehensive cross-evaluation of the set’s performance across different scenarios.
$Q_{s}\subset M_{i},\quad T^{\prime}\subset T_{full}$ (7)
where $Q_{s}$ represents the query trajectories from a specific scene type $s$. The establishment of this baseline metric is crucial, as it supports our hypothesis that trajectory sets sampled from $T_{full}$ will inherently exhibit higher displacement errors compared to scene-specific sets. This phenomenon can be attributed to the inherent diversity and potential noise in the full distribution, whereas scene-specific sets benefit from more focused and contextually relevant trajectory patterns.
VII Experiments
@@
VII-A Evaluation of sets on $T_{full}$
TABLE I: Comparison of coverage-based approach vs. random sampling from the training distribution. ()($\downarrow$)( ↓ ) indicates lower is better, while the inverse holds true to ()($\uparrow$)( ↑ ). The notation 0.2m with Ran-Samp represents the distance threshold applied when calculating metrics for randomly sampled trajectories
@@ VII-A Evaluation of sets on $T_{full}$
Validating the effectiveness of the sets is crucial to comprehend the influence of two primary parameters which govern diversity and plausibility. We employ the metric-driven method [12] and a random sampling to demonstrate the efficacy of our proposed scene-specific set generation concept. Similar to the Covernet [3]; metric-driven method also employs a minADE driven bagging algorithm to reduce each of the clustered sets, we have acquired from the Algorithm 1. Random sampling (Ran-Samp) is used as a baseline, and we report the lowest value as it exhibits a variance of ±$\pm\text{2\%}$%. By comparing the performance of our method against the baseline and analyzing the lower bounds, we gain insights into how effective our set construction is in capturing diverse and plausible trajectories.
Table I illustrates the effect of metric-driven sets when compared with randomly sampled sets generated from the training distribution. As the set size increases, the LB of the metrics linearly decreases, this is attributed to the presence of more trajectories which lie closer to a given $q$, thereby providing good coverage. Sets generated through the metric-driven method perform well as expected, since the cost function enforces diversity between the trajectories in the set when compared with Ran-Samp. On the other hand, from Fig. 3 emerges a wholly different picture. The normalized rate of surviving trajectories, that which are compliant with the scene, is substantially higher for the Ran-Samp and amounts to $10\%$ more trajectories irrespective of the set size.
@@
VII-B Evaluation of scene-specific sets
Utilizing Algorithm 1, we generate two distinct clustered sets: one pertaining to intersections and the other to non-intersections. Subsequently, for each clustered set, adhering to their corresponding sampling criteria, we performed extensive analysis, as can be seen from the Table II.
The DAC metric exhibits relative stability across different set sizes for each method and scenario combination. The Ran-Samp method consistently shows higher DAC values compared to metric-driven, suggesting that it is more effective in generating a set where a larger proportion of the trajectories remain plausible. When comparing the two scenario types, it is evident that both methods yield poorer performance in intersection scenarios compared to non-intersections, as measured between corresponding LB.minADE metrics. This outcome is anticipated due to the increased complexity and diversity of potential trajectories at intersections. However, the performance degradation from non-intersection to intersection scenarios is less evident for Ran-Samp than for metric-driven. As set sizes increase, the displacement error narrows, and the number of surviving trajectories experiences a significant increase, a trend that is also observable in the intersection sets.
We also created a new subsampling algorithm, RIDS or Recursive In-Distribution Subsampling, where we remove overlapping trajectories from a set constructed using Ran-samp. If two trajectories share proximity to each other and lie within a threshold of 0.2m to each other, we consider this trajectory redundant and arbitrarily remove one of them. This leaves us with a set which is underweighted; to counter this, we perform subsampling from the training distribution and recursively perform this step of removing and adding trajectories until the desired $T^{\prime}$ is reached.
@@ VII-B Evaluation of scene-specific sets
TABLE II: Quantitative comparison of Scene-specific sets with different methods on Argoverse 2 training distribution
@@
Method Set Size LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ $T_{full}$ ()($\downarrow$)( ↓ ) DAC()($\uparrow$)( ↑ ) LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ $T_{full}$ ()($\downarrow$)( ↓ ) DAC()($\uparrow$)( ↑ )
1k 0.592 0.670 0.694 0.722 0.735 0.518
1.5k 0.531 0.628 0.673 0.667 0.672 0.500
2k 0.500 0.593 0.677 0.621 0.639 0.490
Metric Driven 2.5k 0.472 0.567 0.664 0.599 0.608 0.484
3k 0.443 0.544 0.704 0.575 0.585 0.477
4k 0.416 0.527 0.692 0.547 0.553 0.468
5k 0.388 0.491 0.687 0.510 0.528 0.469
1k 0.727 0.830 0.827 0.874 0.892 0.691
1.5k 0.673 0.748 0.831 0.789 0.813 0.664
2k 0.601 0.685 0.823 0.729 0.746 0.663
Ran-Samp 2.5k 0.568 0.654 0.842 0.690 0.709 0.678
3k 0.549 0.630 0.833 0.659 0.688 0.692
4k 0.498 0.597 0.835 0.606 0.624 0.681
5k 0.453 0.552 0.863 0.579 0.587 0.679
@@
VIII Results and Discussions
The ultimate goal of this work is to find the answer to the following question: What is the best conceptual framework that can maximize the probability of finding trajectories closer to the ground truth? This led us to investigate and propose a scene-specific set generation method, which, when viewed through the lenses of displacement errors and DAC, gives better performance compared to an uninformed set. The next step is to find the sampling method; From Fig. 3, random sampling provides preliminary evidence, as it retains more trajectories than metric-driven when one wants to achieve a dense representation of the set.
@@ VIII Results and Discussions
Although the metric-driven method was able to provide guarantees of probabilistic completeness to maximize coverage, it did not capture the true distribution, providing insufficient exploration. On the other hand, sampling from a uniform distribution; RIDS was able to retain more trajectories and outperform the metric-driven method. The property of an efficient estimator is that it has the smallest possible variance, indicating that it is closest to the true value of the parameter being estimated. This reasoning that RIDS being an efficient estimator and functions effectively because in certain optimization situations, sampling from the underlying distribution can be more efficient than performing sampling based on some metrics.
Our experiments have shown that the scene-specific set generation method allows us to improve diversity, and more specifically, RIDS; a variant of random sampling works relatively well when considering the environment, resulting in an improved DAC score. The survival rate for the trajectories for sets constructed using Ran-Samp and RIDS outperforms metric-driven methods, given that the minimal gain in displacement errors from metric-driven overshadows the presence of more plausible trajectories. A visual distinction is notable for the non-intersection set in Fig. 1(a), which confirms our prior, as most of the trajectories do not represent significant yaw differences (going straight). The DAC score for the non-intersection with Ran-Samp depicts that cluster-based trajectory sampling works better than expected, since it can retain 86% of the trajectories for $T^{\prime}_{5k}$. This high score indicates that the set can handle the complex geometry of the real-world and is perfectly suited for the intended application. As we said at the start, this showcases that Scene-Specific Set generation can produce sets which can be tuned to represent the ground truth closer to real-world scenes than using a single set to represent diverse situations. Looking at Ran-Samp set created with a threshold rate of 0.2 from Table II; almost one could say that this imbalance between the set size is the reason for the improved DAC value. To counteract this argument, we performed ablation studies as can be seen from Table III and Table IV to confirm that these improved DAC values remain improved when performing the RIDS sampling process for different threshold values.
TABLE III: Ran-Samp @ threshold of 0.5
Set size After Threshold Thresholded Set’s LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ $T_{full}$ ()($\downarrow$)( ↓ ) DAC ()($\uparrow$)( ↑ ) RIDS (LB. minADE) ()($\downarrow$)( ↓ ) RIDS (DAC) ()($\uparrow$)( ↑ )
1k 713 0.807 0.780 0.645 0.747 0.641
1.5k 993 0.770 0.710 0.623 0.694 0.619
2k 1268 0.704 0.661 0.608 0.615 0.617
2.5k 1496 0.659 0.617 0.616 0.585 0.600
3k 1684 0.647 0.598 0.608 0.564 0.591
4k 2204 0.622 0.559 0.590 0.526 0.576
5k 2645 0.597 0.532 0.582 0.497 0.562
TABLE IV: Ran-Samp @ threshold of 0.2
Set size After Threshold Thresholded Set’s LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ $T_{full}$ ()($\downarrow$)( ↓ ) DAC ()($\uparrow$)( ↑ ) RIDS (LB. minADE) ()($\downarrow$)( ↓ ) RIDS (DAC) ()($\uparrow$)( ↑ )
1k 854 0.782 0.780 0.673 0.752 0.676
1.5k 1246 0.737 0.710 0.662 0.712 0.662
2k 1642 0.666 0.661 0.664 0.635 0.680
2.5k 2067 0.621 0.617 0.678 0.599 0.676
3k 2450 0.603 0.598 0.682 0.572 0.671
4k 3237 0.564 0.559 0.671 0.534 0.670
5k 4023 0.538 0.532 0.675 0.503 0.663
IX Conclusion